If a and B are the quadratic polynomial f(x)=xsquare-px+q ,then find the value of a square +B squareand 1/a+1/B

Grade:10

1 Answers

P S S C ASLESH
2 years ago
a and B are the zeroes of the polynomial f(x)=x²-px+q
a+B= -(-p/1)=p
aB=q/1=q
a²+B²=(a+B)²-2aB=p²-2q
1/a+1/B=B+a/ab=p/q